An airplane leaves Los Angeles for Denver at a speed of 430 mph. Thirty minutes later, a plane going from Denver to Los Angeles leaves Denver, which is 850 miles from Los Angeles, at a speed of 500 mph. When they meet, how far are they from Denver?
Answer by mananth(16949)

Plane to Denver 430 km/h
Plane from Denver 500 km/h, 0.5 hours late
Distance 850 miles ,0.5 hours
Plane to Denver has traveled 215 miles before Plane from Denver starts
So distance = 635 when second plane starts
Let time they take to meet be x hours
Time taken to meet is same
t=d/r
combined speed= 930 mph

930x=635
x= 0.68 hours
Plane to Denver 430x +215 m = 508.6 miles
Plane from Denver (850-508.6)= 341.4 miles